# Active Matter as a framework for living systems-inspired Robophysics

### v1

@@ III.6 Testing robot swarms
To test and refine swarm technologies, small-scale and inexpensive testbeds are crucialMokhtarian et al. (2025). As noted by M. Kegeleirs et al.Kegeleirs and Birattari (2025), one of the major obstacles to real-world swarm deployment is the lack of modern, affordable, and reliable experimental platforms. A number of swarm-specific testbeds have been developed, including E-Puck Mondada et al. (2009), Kilobot Rubenstein et al. (2012), Duckietown Paull et al. (2017), and Crazyflie Giernacki et al. (2017). These systems provide standardized, cost-effective environments for studying collective behavior at small scales.
@@
Nevertheless, a common alternative is to build an in-house platform with custom-built García-Pérez et al. (2023) or commercial robots Mañas Álvarez et al. (2023). Among these, the Georgia Tech Robotarium Wilson et al. (2020) stands out as one of the most advanced and influential large-scale experimental facilities. Figure 2 shows a group of five robots. Each robot has an Arduino processor that controls the speed of the motors, receiving feedback signals from the wheel encoders. The Arduino communicates with a Raspberry Pi via a serial port. This Raspberry Pi receives the position or orientation of the robot relative to the arena robot from a server, thanks to an ArUCo-type marker. Some of the robots have battery sensors, and others have infrared distance sensors.
Figure 2: Group of five differential robots from the Robotarium-UCM.
Finally, efficient energy and resource management are critical for the long-term operation of swarm systems. Developing strategies to monitor and balance power consumption across agents will be essential for scaling swarm robotics toward real-world, autonomous applications.
IV Active Matter as an effective framework for Robophysics
@@
Intelligent Soft Matter has been recently developed as a field of research aimed at creating biological-inspired materials with life-like capabilities, such as perception, learning, memory, and adaptive behaviorBaulin et al. (2025). Active matter has emerged in recent years as a robust framework for studying collective behaviors in both livingMarchetti et al. (2013); Bechinger et al. (2016); Gompper et al. (2025) and syntheticJu et al. (2025) systems. It focuses on ensembles of self-propelled agents, whether biological (e.g., bacteria, fish) or artificial (e.g., colloids), that consume energy to generate motion and stresses, thereby driving the system out of thermodynamic equilibrium Marchetti et al. (2013). These agents can exhibit striking emergent phenomena, such as flocking or swarming, that arise purely from local interactions. Synthetic micro/nanorobots could be have potential applications in biomedicine and environmental remediation.Ju et al. (2025) However, in contrast to biological systems, most active matter models lack internal decision-making and goal-oriented dynamics: while they capture emergent motion, they rarely encode the shared purposes that living collectives achieve, such as predator evasion, energy optimization, or efficient foraging Te Vrugt and Wittkowski (2025).
@@
Recent developments highlight ongoing efforts to close this gap. New models seek to incorporate perception, adaptability, and purposeful behavior into synthetic active systems March-Pons et al. (2025, 2024); Acosta-Soba et al. (2025); Abdoli et al. (2025); Löwen and Liebchen (2025); Cai et al. (2025). At the same time, machine learning methods have been nowadays increasingly applied to active matter systemsBapst et al. (2020); Boattini et al. (2021); Clegg (2021); Wetzel et al. (2025); Jung et al. (2025); Janzen et al. (2024); Cichos et al. (2020); Janzen et al. (2023); Volpe et al. (2024). In particular, neural networks have been integrated into numerical simulations and trained on experimental data, enabling the simulations to capture complex, data-driven cell behaviors Minartz et al. (2025); neural network classifiers have been trained on numerical simulation data to distinguish motile from non-motile cells Braat et al. (2025); and graph neural networks have been used to extract the underlying pairwise interaction rules governing the dynamics of self-propelled colloids from experimental data of active colloidsRuiz-Garcia et al. (2024).
@@
Moreover, beyond analysis, machine learning methods have been applied for inverse designLiu et al. (2025), which seeks to identify microscopic ingredients or control parameters leading to a desired macroscopic response. Reinforcement learningSutton et al. (1998); Brunton and Kutz (2022) (RL) is particularly well suited for this task, as it mimics trial-and-error strategies typical of biological systems. In Reinforcement Learning, an agent interacts with its environment, receiving positive or negative feedback that guides it toward actions maximizing a long-term reward Brunton and Kutz (2022). Such approaches have already been applied to soft and active matter systems, for example, to discover optimal protocols for steering active particles between states at minimal energetic cost Falk et al. (2021), to dynamically control the self-assembly of quasicrystalline structures Lieu and Yoshinaga (2025), and to learn control strategies for non-equilibrium dynamics using only local, imperfect gradient information Floyd et al. (2024).
V Outlook
@@
Looking forward, we believe that active matter physics can be considered as the framework to tackle Robophysics. Active Matter allows us to extract the minimal ingredients needed to reproduce a Nature-inspired collective motion, such as purpose, perception, communication, and adaptive feedback, and to translate them into robotic swarms. Since machine learning and reinforcement learning have recently proven to be powerful tools for analyzing and controlling complex dynamical systems, their integration with active matter offers a path toward a conceptual framework that identifies the essential ingredients of living collectives and enables their implementation in robotic systems. Achieving this goal will require an interdisciplinary approach: physics to identify the governing principles, robotics and engineering to implement them in hardware, and computer science to provide the communication and data-processing infrastructure. By merging these perspectives, robophysics will move beyond locomotion toward programmable collectives that approach the functionality of living systems.
VI Acknowledgments
@@
C.V. acknowledges fundings IHRC22/00002 and Proyecto PID2022-140407NB-C21 funded by MCIN/AEI /10.13039/501100011033 and FEDER, UE. DMF and GJ acknowledges support from MINECO (MCIN/AEI Grant No. PID2023-148991NA-I00). DMF also is funded through the Ramon y Cajal Program of the Spanish Ministry of Science and Innovation (Grant No. RYC2023-043002-I). J.J. and L.G-P. acknowledge support from iRoboCity2030-CM, Robotica Inteligente (TEC2024/TEC62TEC-2024/TEC-62), funded by the Programas de Actividades $I+D$ en Tecnologias en la Comunidad de Madrid. G.M acknowledges fundings Sapienza Award Horizon Europe - 2024.

### v2

@@ 3.6 Testing robot swarms
To test and refine swarm technologies, small-scale and inexpensive testbeds are crucial[67]. As noted by M. Kegeleirs et al.[54], one of the major obstacles to real-world swarm deployment is the lack of modern, affordable, and reliable experimental platforms. A number of swarm-specific testbeds have been developed, including E-Puck [68], Kilobot [76], Duckietown [70], and Crazyflie [42]. These systems provide standardized, cost-effective environments for studying collective behavior at small scales.
@@
Nevertheless, a common alternative is to build an in-house platform with custom-built [40] or commercial robots [60], such as those built at the Georgia Tech Robotarium [93] or at UCM-Robotarium. Figure 2 shows a group of five robots (UCM-Robotarium), each with an Arduino processor controlling the motors’ speed, receiving feedback signals from the wheel encoders.
FIG. 2: Group of five differential robots from the Robotarium-UCM. The Arduino communicates with a Raspberry Pi via a serial port. This Raspberry Pi receives the position or orientation of the robot relative to the arena robot from a server, thanks to an ArUCo-type marker. Some of the robots have battery sensors, and others have infrared distance sensors.
Finally, efficient energy and resource management are critical for the long-term operation of swarm systems. Developing strategies to monitor and balance power consumption across agents is essential for scaling swarm robotics toward real-world, autonomous applications. These challenges can be naturally organized according to their functional role and level of dependency. At a foundational level, the availability of experimental platforms, scalability and decentralization, the cooperative localization in GPS-deprived environments, the coupling between communication and motion, and the energy/resources management are essential prerequisites for the operation of large robotic swarms. Building upon this physical and operational substrate, higher-level challenges are the emergence of coordinated collective behaviors, including a shared purpose and goal-oriented dynamics. Finally, adaptive and resilience-related aspects such as learning, robustness, and fault tolerance represent advanced capabilities that can be progressively developed once a coherent and scalable collective system is in place.
Notably, the foundational and coordination-level challenges are precisely those for which the Robophysics and active matter frameworks provide particularly powerful conceptual and experimental tools, as they explicitly address the coupling between physical interactions, motion, communication, and emergent collective dynamics.
4 Active Matter as an effective framework for Robophysics
@@
Intelligent Soft Matter has been recently developed as a field of research aimed at creating biological-inspired materials with life-like capabilities, such as perception, learning, memory, and adaptive behavior[10]. Active matter has emerged in recent years as a robust framework for studying collective behaviors in both living[63, 11, 45] and synthetic[52] systems. It focuses on ensembles of self-propelled agents, whether biological (e.g., bacteria, fish) or artificial (e.g., colloids), consuming energy to generate motion and stresses, thereby driving the system out of thermodynamic equilibrium [63]. These agents can exhibit striking emergent phenomena, such as flocking or swarming, purely arising from local interactions. Synthetic micro/nanorobots could have potential applications in biomedicine and environmental remediation.[52] Flocking emerge in suspensions of self-propelled Janus colloids with anisotropic repulsioin at the front/rear characterised by turn-away torques [30]. Light-driven synthetic micro-swimmers form dense swarms capable of trapping and transporting passive loads [12], whereas dense binary mixtures of active Janus colloids (with distinct motilities and independently tunable alignment [5]) aggregate into self-organized and highly dynamic polar clusters. In contrast to biological systems, most active matter models lack internal decision-making and goal-oriented dynamics: while capturing emergent motion, they rarely encode shared purposes that living collectives achieve, such as predator evasion, energy optimization, or efficient foraging [85]. Recent developments highlight ongoing efforts to close this gap. New models seek to incorporate perception, adaptability, and purposeful behavior into synthetic active systems [62, 61, 2, 1, 58, 22]. Cooperative transport spontaneously emerges in swarms of stochastic self-propelled robots [7] and in swarms of robots[23], with applications in embodied decentralized control. Swarms of active particles confined in a flexible scaffold have been proposed to build functional soft structures. Mindless rod-like motile robots in confinement undergo collective behavior near boundaries[31], and are capable to perform simple tasks such as pulling a load [16].
@@
At the same time, machine learning methods have been nowadays increasingly applied to active matter systems[8, 15, 26, 92, 53, 50, 25, 49, 87]. In particular, neural networks have been integrated into numerical simulations and trained on experimental data, enabling simulations to capture complex, data-driven cell behaviors [65]. Neural network classifiers have been trained on numerical data to distinguish motile from non-motile cells [17], and graph neural networks have been used to extract the underlying pairwise interaction governing the dynamics of active colloids from experimental data [77].
@@
Beyond analysis, machine learning methods have been applied for inverse design[57], seeking to identify microscopic ingredients or control parameters leading to a desired macroscopic response. Reinforcement learning[81, 20] (RL) is particularly suited for this task, as it mimics trial-and-error strategies typical of biological systems. In Reinforcement Learning, an agent interacts with its environment, receiving positive or negative feedback that guides it toward actions maximizing a long-term reward [20]. Such approaches have already been applied to soft and active matter systems to discover optimal protocols for steering active particles between states at minimal energetic cost [36], to dynamically control self-assembly of quasicrystalline structures [56] and to learn control strategies for non-equilibrium dynamics using only local gradient information [38]. Recent work has also brought RL and related adaptive feedback schemes into a physical setting of mechanically interacting robotic swarms, demonstrating adaptive phototaxis via positive and negative feedback self-alignment [66] and morphological computation with decentralized learning in sterically interacting robots [95].
5 Outlook
Looking forward, we believe that active matter physics can be considered as the framework to tackle Robophysics. Active Matter allows to extract the minimal ingredients needed to reproduce a Nature-inspired collective motion including purpose, perception, communication, and adaptive feedback, and to translate all into robotic swarms.
@@ 5 Outlook
Since machine learning and reinforcement learning have recently proven to be powerful tools for analyzing and controlling complex dynamical systems, their integration with active matter offers a path toward a conceptual framework that identifies the essential ingredients of living collectives and enables their implementation in robotic systems. Achieving this goal will require an interdisciplinary approach: physics to identify the governing principles, robotics and engineering to implement them in hardware, and computer science to provide the communication and data-processing infrastructure. By merging these perspectives, active matter will leverage robophysics to move beyond locomotion toward programmable collectives that approach the functionality of living systems.
6 Acknowledgments
@@
C.V. acknowledges fundings IHRC22/00002 and Proyecto PID2022-140407NB-C21 funded by MCIN/AEI /10.13039/501100011033 and FEDER, UE. DMF and GJ acknowledge support from MINECO (MCIN/AEI Grant No. PID2023-148991NA-I00). DMF also is funded through the Ramon y Cajal Program of the Spanish Ministry of Science and Innovation (Grant No. RYC2023-043002-I). J.J. and L.G-P. acknowledge support from iRoboCity2030-CM, Robotica Inteligente (TEC2024/TEC62TEC-2024/TEC-62), funded by the Programas de Actividades $I+D$ en Tecnologias en la Comunidad de Madrid. G.M acknowledges fundings Sapienza Award Horizon Europe - 2024.